Use the quadratic formula to solve each of the following quadratic equations.
2x^2-5x=3
thank you

Answer by checkley75(3666)

2X^2-5X-3=0
X=(5+-SQRT[-5^2-4*2*-3])/2*2
X=(5+-SQRT[25+24])/4
X=(5+-SQRT49)/4
X=(5+-7)/4
X=(5+7)/4
X=12/4
X=3 ANSWER
X=(5-7)/4
X=-2/4
X=-.5 ANSWER.